OpenStudy (anonymous):

Part 1: Use the quadratic formula to solve x2 + 8x = –2. (3 points) Part 2: Using complete sentences, explain the process you used. (2 points) Part 3: Why is the quadratic formula the best method to use?

OpenStudy (paxpolaris):

\[x^2+8x=-2\]\[\implies \large x^2+8x+2=0\] \[when\ ax^2+bx+c=0,\ the\ quadratic\ formula\ is:\]\[\Large x= {-b \pm \sqrt{b^2-4ac} \over 2a}\]